OpenStudy (anonymous):

jimmy ran 3 miles from home , then turned north and jogged 4 miles , how far was jimmy from home

OpenStudy (anonymous):

r you working with pythagorean theorem

OpenStudy (anonymous):

I honestly don't know

OpenStudy (anonymous):

can you help ?

OpenStudy (anonymous):

um... the answer would be 5 miles if you are working with a pythagorean theorem. because it is a 3,4,5 triangle

OpenStudy (yanasidlinskiy):

\[c ^{2}=a ^{2}+b ^{2}\]

OpenStudy (yanasidlinskiy):

\[c^2=3^2+4^2\]

OpenStudy (yanasidlinskiy):

\[c^2=9+16=c^2=25\]

OpenStudy (yanasidlinskiy):

\[c=\sqrt{25}\]

OpenStudy (yanasidlinskiy):

c=5